OpenStudy (anonymous):

can someone check this out? the product of 7 and of a # is 12 greater than 5 times the #, what is the #?

OpenStudy (anonymous):

f(2)=3(2)^3 - 4(3)^2 + 2 -1 f(2)=3(8) - 4(9) + 1 f(2)=24 - 36 + 1 Is this what you got so far when you were working out the problem?

OpenStudy (anonymous):

\[3x^3 - 4x^2 + x - 1 i get this : 3 (2)^3 - 4 (2)^2 + 2 -1\]

OpenStudy (anonymous):

okay, pretty good so far.....

OpenStudy (anonymous):

and after that I get , 3 (8) - 4 (4) + 2 - 1 and after I solve it I get 3 as my answer

OpenStudy (anonymous):

24 - 16 + 2 - 1 -------------- 8 + 2 - 1 -------------- 9 How did you get 3?

OpenStudy (anonymous):

Because the answer should be 9...

OpenStudy (anonymous):

just saw that by mistake i wrote down the wrong numbers.. but thanks could you help me with the other question. with that one i have no idea how to solve it

OpenStudy (anonymous):

I am sorry I can't because I have to go, but I will tag someone who can help you, he's REALLY good at math....

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

PLEASE HELP @SolomonZelman

OpenStudy (anonymous):

thanks for your help tho

OpenStudy (anonymous):

np ;p

OpenStudy (anonymous):

@phi @amistre64

OpenStudy (anonymous):

PLEASE HELP HER